# The non-Abelian geometry, topology, and dynamics of a nonreciprocal Su-Schrieffer-Heeger ladder

### v1

@@ Appendix C A gauge-invariant winding number for non-Hermitian multi-band systems with chiral symmetries
$U_{k}^{-1}H_{k}U_{k}=\left(\begin{array}[]{cc}E_{k}&0\\ 0&-E_{k}\end{array}\right).$ (111)
where $U_{k}^{-1}=(U_{+}^{(l)},U_{-}^{(l)})^{\dagger}$ and $E_{k}\equiv\text{diag}(\epsilon_{k1},\cdots,\epsilon_{kN})$. Thus, we can rewrite $H_{k}$ using the eigenvectors and the eigenenergies as
@@
To define the topology, one can construct a $Q$-matrix [63] that has identical eigenvectors but with eigenvalues ±$\pm E_{k}$ of the eigenvectors in $\mathbb{G}_{\pm}$ subspaces collapsed to ±$\pm 1$, respectively:
@@
$Q_{k}^{(b)}\equiv U_{s}^{{\dagger}}Q_{k}U_{s}\equiv\left(\begin{array}[]{cc}0&q_{k}\\ q^{-1}_{k}&0\end{array}\right)$ (118)
where $q_{k}$ is an $N\times N$ invertible matrix (i.e., $\det q_{k}\neq 0$). That the two block matrices are invertible to each other is derived from $[Q_{k}^{(b)}]^{2}=Q_{k}^{2}=I_{2N}$. Then, we can define the winding number:
@@
$=$ d $\text{d\,tr}\ln\Lambda=\text{d}\ln\det\Lambda=\text{d}\ln\det U\Lambda U^{-1}$
@@
where we use the eigenvalue decomposition $U^{-1}q_{k}U=\Lambda\equiv\text{diag}(\{\lambda_{s}\})$ with $\lambda_{s}\,(s=1,\cdots,N)$ being the eigenvalues, and $\ln\Lambda\equiv\text{diag}(\{\ln\lambda_{s}\})$. This winding number $w$ defines a map from $k\in S^{1}$ onto an $N\times N$ invertible matrix $q_{k}\in GL(N,\mathbb{C})$ to reflect the topology of the bipartition of the Hilbert space into two subspaces $\mathbb{G}_{\pm}$ in BZ. This map is a fundamental homotopy group $\pi_{1}(GL(N,\mathbb{C}))\in\mathbb{Z}$. Specifically, when $H_{k}$ and thus $Q_{k}$ and $Q_{k}^{(b)}$ are Hermitian, the invertible matrix $q_{k}$ becomes unitary (i.e., $q_{k}^{\dagger}=q_{k}^{-1}$), reducing the fundamental homotopy group from $\pi_{1}(GL(N,\mathbb{C}))$ to $\pi_{1}(U(N))\in\mathbb{Z}$.
@@
$I_{N}$ $=$ $[U_{-}^{(l){\dagger}}U_{s}][U_{s}^{\dagger}U_{-}^{(r)}]=U_{1}^{(l){\dagger}}U_{1}^{(r)}+U_{2}^{(l){\dagger}}U_{2}^{(r)}$
00 0 $=$ $[U_{+}^{(l){\dagger}}U_{s}][U_{s}^{\dagger}U_{-}^{(r)}]=U_{1}^{(l){\dagger}}U_{1}^{(r)}-U_{2}^{(l){\dagger}}U_{2}^{(r)}$
@@
$=$ $w_{-}+\frac{1}{\pi i}\int_{k\in\text{BZ}}\text{tr}\,R_{k}^{-1}\text{d}R_{k}$
$=$ $w_{-}+\frac{1}{\pi i}\int_{k\in\text{BZ}}\text{tr}\,\text{d}\ln\Lambda$
@@
$=$ $w_{-}+\frac{1}{\pi}\sum_{n=1}^{N}\int_{k\in\text{BZ}}\text{d}\varphi_{n}$
$=$ $w_{-}+2\times\text{Integer},$
where we use the same trick as in Eq. (120), $U^{-1}R_{k}U=\Lambda\equiv\text{diag}(\{r_{n}e^{i\varphi_{n}}\})~{}(n=1,\cdots,N)$ with $r_{n}\geq 0$ and $\varphi_{n}\in\mathbb{R}$, and notice that $r_{n}$ is periodic in $k$ but $\varphi_{n}$ is periodic modulo $2\pi$, since $R_{k}$ is a periodic function of $k$. Eq. (136) demonstrates that this definition of winding number is only valid modulo $2$, i.e., $w_{-}\in\mathbb{Z}_{2}$, much coarser than $w\in\mathbb{Z}$. It is interesting to note that the two kinds of winding numbers come from different parts of the matrix of the non-Hermitian non-Abelian Berry connection for the whole bands:
@@
The following lists more possibly useful relations for understanding the properties of this winding number:
$\frac{\det h_{k}^{(+)}}{\det h_{k}^{(-)}}=(\det q_{k})^{2},~{}\det h_{k}^{(+)}\det h_{k}^{(-)}=(\det E_{k})^{2}.$ (146)
Appendix D The relation between the dynamics and the Wilson line for non-Hermitian systems
When we add an external constant force $F$ along, say, the right direction of the ladder, the time-dependent Schrödinger equation becomes ($\hbar=1$) [7]
$i{\partial_{t}}|\psi(t)\rangle=(\hat{H}_{\text{HN}}-F\hat{X})|\psi(t)\rangle,$ (147)
where $\partial_{t}\equiv\partial/\partial t$ and for simplicity the position operator $\hat{X}$ is defined with respect to the positions of unit cells:
$\hat{X}:\hat{a}_{j\sigma}^{({\dagger})}\rightarrow j\hat{a}_{j\sigma}^{({\dagger})},\,\hat{b}_{j\sigma}^{({\dagger})}\rightarrow j\hat{b}_{j\sigma}^{({\dagger})}$ (148)
in real space, and correspondingly
$\hat{X}:\hat{\psi}_{k}^{{\dagger}}\rightarrow-i\partial_{k}\hat{\psi}_{k}^{{\dagger}},\,\hat{\psi}_{k}\rightarrow i\partial_{k}\hat{\psi}_{k},$ (149)
in $k$-space with $\partial_{k}\equiv\partial/\partial_{k}$. We rewrite the Hamiltonian (49) in the diagonal form as
$\hat{H}_{\text{HN}}=\sum_{k\in{\text{B}Z}}\hat{u}_{k}^{(r){\dagger}}\Lambda_{k}\hat{u}^{(l)}_{k}=\sum_{k\in{\text{B}Z}}\sum_{n=1}^{4}\epsilon_{kn}\hat{u}_{kn}^{(r){\dagger}}\hat{u}^{(l)}_{kn},$ (150)
where $\hat{u}_{k}^{(r){\dagger}}\equiv(\hat{u}_{k1}^{(r){\dagger}},\hat{u}_{k2}^{(r){\dagger}},\hat{u}_{k3}^{(r){\dagger}},\hat{u}_{k4}^{(r){\dagger}})=\hat{\psi}_{k}^{\dagger}U_{k}$, $\hat{u}_{k}^{(l)}\equiv(\hat{u}_{k1}^{(l)},\hat{u}_{k2}^{(l)},\hat{u}_{k3}^{(l)},\hat{u}_{k4}^{(l)})^{T}=U_{k}^{-1}\hat{\psi}_{k}$, and $\Lambda_{k}=\text{diag}(\epsilon_{k1},\epsilon_{k2},\epsilon_{k3},\epsilon_{k4})$. We set the initial state as a superposition of the Bloch states at $k_{0}$, i.e., $|\psi(0)\rangle=\hat{u}_{k}^{(r){\dagger}}\alpha(0)|0\rangle$, where the coefficient vector $\alpha(t)=(\alpha_{1}(t),\alpha_{2}(t),\alpha_{3}(t),\alpha_{4}(t))^{T}$ with $\alpha_{n}(t)$ being the amplitude of the Bloch state in the $n$th band at time $t$, and we normalize the coefficients at $t=0$ by $\sum_{n}|\alpha_{n}(0)|^{2}=1$. With the ansatz that the initial state uniformly evolves to another superposition state $|\psi(t)\rangle=\hat{u}_{k(t)}^{\dagger}\alpha(t)|0\rangle$ at $k(t)=k_{0}+Ft$, using Eq. (147) we have the following equations of motion:
@@ Appendix D The relation between the dynamics and the Wilson line for non-Hermitian systems
is just the non-Hermitian non-Abelian Berry connection Eq. (137) for the whole bands. The solution to Eq. (151) is
[…]
@@
$\approx$ $e^{-i\Lambda_{k}dk/F}(I-iA_{k}dk)$
$=$ $e^{-i\Lambda_{k}dk/F}(I-U_{k}^{-1}dU_{k})$
@@
$V_{k_{0}\rightarrow k_{f}}$ $=$ $e^{-i\Lambda_{k_{f}}dk/F}U_{k_{f}}^{-1}{U}_{k_{l-1}}\cdots e^{-i\Lambda_{k_{i}}dk/F}{U}_{{k}_{i}}^{-1}{U}_{{k}_{i-1}}$ (155)
$\cdots e^{-i\Lambda_{k_{1}}dk/F}{U}_{{k}_{1}}^{-1}{U}_{{k}_{0}}.$
where $(k_{1},\cdots,k_{l-1})$ are the $l-1$ joint points of the $l$ pieces of equal intervals along the path.
@@
[…]
@@
$W_{k_{0}\rightarrow k_{f}}=U_{k_{f}}^{-1}{U}_{k_{l-1}}\cdots{U}_{{k}_{i}}^{-1}{U}_{{k}_{i-1}}\cdots{U}_{{k}_{1}}^{-1}{U}_{{k}_{0}}=U_{k}^{-1}U_{k_{0}}.$ (157)
is a four-band Wilson line for non-Hermitian systems along the path from $k_{0}$ to $k_{f}$ in $k$-space [50], reduced from $V_{k_{0}\rightarrow k_{f}}$ in Eq. (155). This Wilson line defined using the biorthonormal basis is a non-Hermitian generalization of the Wilson line in Hermitian systems [7, 9]. Apparently, $W_{k_{f}\rightarrow k_{0}}=W_{k_{0}\rightarrow k_{f}}^{-1}$. If $k_{f}=k_{0}+2\pi n$ ($n\in$ Integer), the Wilson line becomes a Wilson loop due to the periodicity of the BZ, and thus, $W_{k_{0}\rightarrow k_{0}+2\pi n}=I$, indicating the state returns to the initial state, i.e., $|\psi(t)\rangle=|\psi(0)\rangle$.
@@
where (±)plus-or-minus($\pm$)( ± ) means the upper/lower $2\times 2$ diagonal blocks in the labeled matrices and $2\times 1$ blocks in labeled column vectors, and $W^{(\pm)}_{k_{0}\rightarrow k_{f}}$ are the two-band Wilson lines for non-Hermitian systems in corresponding subspaces. Likewise, $W^{(\pm)}_{k_{0}\rightarrow k_{f}}$ can also be numerically calculated as follows:
$W^{(\pm)}_{k_{0}\rightarrow k}$ $=$ $U_{\pm,k}^{(l){\dagger}}{U}_{\pm,k_{l-1}}^{(r)}\cdots{U}_{\pm,k_{i}}^{(l){\dagger}}{U}_{\pm,{k}_{i-1}}^{(r)}\cdots{U}_{\pm,{k}_{1}}^{(l){\dagger}}{U}_{\pm,{k}_{0}}^{(r)}$ (159)
$=$ $U_{\pm,k}^{(l){\dagger}}P_{\pm,k_{l-1}}\cdots P_{\pm,{k}_{i}}\cdots P_{\pm,{k}_{1}}U_{\pm,{k}_{0}}^{(r)},$

### v2

@@ Appendix C A gauge-invariant winding number for non-Hermitian multi-band systems under the chiral symmetry
$U_{k}^{-1}H_{k}U_{k}=\left(\begin{array}[]{cc}E_{k}&0\\ 0&-E_{k}\end{array}\right),$ (111)
where $U_{k}^{-1}=(U_{+}^{(l)},U_{-}^{(l)})^{\dagger}$ and $E_{k}\equiv\text{diag}(\epsilon_{k1},\dots,\epsilon_{kN})$. Thus, we can rewrite $H_{k}$ using the eigenvectors and the eigenenergies as
@@
To define the topology, one can construct a $Q$-matrix [67] that has identical eigenvectors but with eigenvalues $\pm E_{k}$ of the eigenvectors in $\mathbb{G}_{\pm}$ subspaces collapsed to $\pm 1$, respectively:
@@
$Q_{k}^{(b)}\equiv U_{s}^{{\dagger}}Q_{k}U_{s}\equiv\left(\begin{array}[]{cc}0&q_{k}\\ q^{-1}_{k}&0\end{array}\right),$ (118)
where $q_{k}$ is an $N\times N$ invertible matrix (i.e., $\det q_{k}\neq 0$). That the two block matrices are invertible to each other is derived from $Q_{k}^{(b)}]^{2}=Q_{k}^{2}=I_{2N}$. Then, we can define the winding number:
@@
$=$ $\text{d\,tr}\ln\Lambda=\text{d}\ln\det\Lambda=\text{d}\ln\det U\Lambda U^{-1}$
@@
where we use the eigenvalue decomposition $U^{-1}q_{k}U=\Lambda\equiv\text{diag}(\{\lambda_{s}\})$ with $\lambda_{s}\,(s=1,\dots,N)$ being the eigenvalues, and $\ln\Lambda\equiv\text{diag}(\{\ln\lambda_{s}\})$. This winding number $w$ defines a map from $k\in S^{1}$ onto an $N\times N$ invertible matrix $q_{k}\in GL(N,\mathbb{C})$ to reflect the topology of the bipartition of the Hilbert space into two subspaces $\mathbb{G}_{\pm}$ in BZ. This map is a fundamental homotopy group $\pi_{1}(GL(N,\mathbb{C}))\in\mathbb{Z}$. Specifically, when $H_{k}$ and thus $Q_{k}$ and $Q_{k}^{(b)}$ are Hermitian, the invertible matrix $q_{k}$ becomes unitary (i.e., $q_{k}^{\dagger}=q_{k}^{-1}$), reducing the fundamental homotopy group from $\pi_{1}(GL(N,\mathbb{C}))$ to $\pi_{1}(U(N))\in\mathbb{Z}$.
@@
$I_{N}$ $=$ $[U_{-}^{(l){\dagger}}U_{s}][U_{s}^{\dagger}U_{-}^{(r)}]=U_{1}^{(l){\dagger}}U_{1}^{(r)}+U_{2}^{(l){\dagger}}U_{2}^{(r)},$
$0$ $=$ $[U_{+}^{(l){\dagger}}U_{s}][U_{s}^{\dagger}U_{-}^{(r)}]=U_{1}^{(l){\dagger}}U_{1}^{(r)}-U_{2}^{(l){\dagger}}U_{2}^{(r)},$
@@
$=$ $w_{-}+\frac{1}{\pi i}\int_{k\in\text{BZ}}\text{tr}\,R_{k}^{-1}\text{d}R_{k}=w_{-}+\frac{1}{\pi i}\int_{k\in\text{BZ}}\text{tr}\,\text{d}\ln\Lambda$
@@
$=$ $w_{-}+\frac{1}{\pi}\sum_{n=1}^{N}\int_{k\in\text{BZ}}\text{d}\varphi_{n}=w_{-}+2\times\text{Integer},$
where we use the same trick as in Eq. (120), $U^{-1}R_{k}U=\Lambda\equiv\text{diag}(\{r_{n}e^{i\varphi_{n}}\})~(n=1,\dots,N)$ with $r_{n}\geq 0$ and $\varphi_{n}\in\mathbb{R}$, and notice that $r_{n}$ is periodic in $k$ but $\varphi_{n}$ is periodic modulo $2\pi$, since $R_{k}$ is a periodic function of $k$. Equation (136) demonstrates that this definition of winding number is only valid modulo $2$, i.e., $w_{-}\in\mathbb{Z}_{2}$, much coarser than $w\in\mathbb{Z}$. It is interesting to note that the two kinds of winding numbers come from different parts of the matrix of the non-Hermitian non-Abelian Berry connection for the whole bands:
@@
Two useful relations are listed here:
$\frac{\det h_{k}^{(+)}}{\det h_{k}^{(-)}}=(\det q_{k})^{2},~\det h_{k}^{(+)}\det h_{k}^{(-)}=(\det E_{k})^{2}$ (146)
for understanding the properties of this winding number.
Appendix D Relation between the dynamics and the Wilson line for non-Hermitian systems
When we add an external constant force $F$ along the ladder, i.e., the Hamiltonian in Eq. (1) becomes $\hat{H}-F\hat{X}$, where for simplicity the position operator $\hat{X}$ is defined only with respect to the positions of unit cells, i.e.,
$\hat{X}=\sum_{j\sigma}j(\hat{a}_{j\sigma}^{{\dagger}}\hat{a}_{j\sigma}+\hat{b}_{j\sigma}^{{\dagger}}\hat{b}_{j\sigma}).$ (147)
Following the similar derivation from Eq. (4) to (5), the short-time dynamics before a quantum jump occurs can be well captured by the time-dependent non-Hermitian Schrödinger equation ($\hbar=1$):
$i{\partial_{t}}|\psi(t)\rangle=(\hat{H}_{\text{NH}}-F\hat{X})|\psi(t)\rangle,$ (148)
where $\partial_{t}\equiv\partial/\partial t$, and the loss term $i\delta\hat{N}$ in Eq. (5), which only contributes an overall decaying factor to the dynamics, is ignored. Here, we rewrite the Hamiltonian (49) in the diagonal form:
$\hat{H}_{\text{NH}}=\sum_{k\in{\text{B}Z}}\hat{u}_{k}^{(r){\dagger}}\Lambda_{k}\hat{u}^{(l)}_{k}=\sum_{k\in{\text{B}Z}}\sum_{n=1}^{4}\epsilon_{kn}\hat{u}_{kn}^{(r){\dagger}}\hat{u}^{(l)}_{kn}$, (149)
where $\hat{u}_{k}^{(r){\dagger}}\equiv[\hat{u}_{k1}^{(r){\dagger}},\hat{u}_{k2}^{(r){\dagger}},\hat{u}_{k3}^{(r){\dagger}},\hat{u}_{k4}^{(r){\dagger}}]=\hat{\psi}_{k}^{\dagger}U_{k}$, $\hat{u}_{k}^{(l)}\equiv[\hat{u}_{k1}^{(l)},\hat{u}_{k2}^{(l)},\hat{u}_{k3}^{(l)},\hat{u}_{k4}^{(l)}]^{T}=U_{k}^{-1}\hat{\psi}_{k}$, and $\Lambda_{k}=\text{diag}(\epsilon_{k1},\epsilon_{k2},\epsilon_{k3},\epsilon_{k4})$. We set the initial state as a superposition of Bloch states of different bands at $k_{0}$, i.e., $|\psi(0)\rangle=\hat{u}_{k}^{(r){\dagger}}\alpha(0)|0\rangle$, where the coefficient vector $\alpha(t)=[\alpha_{1}(t),\alpha_{2}(t),\alpha_{3}(t),\alpha_{4}(t)]^{T}$ with $\alpha_{n}(t)$ being the amplitude of Bloch state in the $n$th band at time $t$, and we normalize the coefficients at $t=0$ with $\sum_{n}|\alpha_{n}(0)|^{2}=1$. With the ansatz that the initial state uniformly evolves to another superposition of Bloch states at $k(t)=k_{0}+Ft$, i.e., $|\psi(t)\rangle\approx\hat{u}_{k(t)}^{(r){\dagger}}\alpha(t)|0\rangle$ [7], Eq. (148) reduces to
@@ Appendix D Relation between the dynamics and the Wilson line for non-Hermitian systems
with $\partial_{k}\equiv\partial/\partial_{k}$ is just the non-Hermitian non-Abelian Berry connection Eq. (137) for the whole bands. The solution to Eq. (150) is
$\alpha(t)$ $=$ $\mathcal{T}\exp\Big{\{}-i\int_{0}^{t}dt\big{[}\Lambda_{k(t)}+FA_{k(t)}\big{]}\Big{\}}\alpha(0)$ (152)
$=$ $\mathcal{P}\exp\Big{\{}-i\int_{k_{0}}^{k_{f}}dk\big{(}\Lambda_{k}/F+A_{k}\big{)}\Big{\}}\alpha(0)$
@@
$\approx$ $e^{-i\Lambda_{k}dk/F}(I-iA_{k}dk)=e^{-i\Lambda_{k}dk/F}(I-U_{k}^{-1}dU_{k})$
@@
$V_{k_{0}\rightarrow k_{f}}$ $=$ $e^{-i\Lambda_{k_{f}}dk/F}U_{k_{f}}^{-1}{U}_{k_{l-1}}\dots e^{-i\Lambda_{k_{i}}dk/F}{U}_{{k}_{i}}^{-1}{U}_{{k}_{i-1}}$ (154)
$\dots e^{-i\Lambda_{k_{1}}dk/F}{U}_{{k}_{1}}^{-1}{U}_{{k}_{0}}.$
where $(k_{1},\dots,k_{l-1})$ are the l1l-1 joint points of the $l$ pieces of equal intervals along the path.
@@
$\alpha(t)\approx\mathcal{P}\exp\Big{[}-i\int_{k_{0}}^{k_{f}}dkA_{k}\Big{]}\alpha(0)\equiv W_{k_{0}\rightarrow k_{f}}\alpha(0),$ (155)
@@
$W_{k_{0}\rightarrow k_{f}}=U_{k_{f}}^{-1}{U}_{k_{l-1}}\dots U_{{k}_{i}}^{-1}{U}_{{k}_{i-1}}\dots U_{{k}_{1}}^{-1}{U}_{{k}_{0}}=U_{k}^{-1}U_{k_{0}}.$ (156)
is a four-band Wilson line for non-Hermitian systems along the path from $k_{0}$ to $k_{f}$ in $k$ space [50], reduced from $V_{k_{0}\rightarrow k_{f}}$ in Eq. (154). This Wilson line defined using the biorthonormal basis is a non-Hermitian generalization of the Wilson line in Hermitian systems [7, 9]. Apparently, $W_{k_{f}\rightarrow k_{0}}=W_{k_{0}\rightarrow k_{f}}^{-1}$. If $k_{f}=k_{0}+2\pi n$ ($n\in$ Integer), the Wilson line becomes a Wilson loop due to the periodicity of the BZ, and thus, $W_{k_{0}\rightarrow k_{0}+2\pi n}=I$, indicating the state returns to the initial state, i.e., $|\psi(t)\rangle=|\psi(0)\rangle$.
@@
where $(\pm)$ means the upper/lower $2\times 2$ diagonal blocks in the labeled matrices and $2\times 1$ blocks in labeled column vectors, and $W^{(\pm)}_{k_{0}\rightarrow k_{f}}$ are the two-band Wilson lines for non-Hermitian systems in corresponding subspaces. Likewise, $W^{(\pm)}_{k_{0}\rightarrow k_{f}}$ can also be numerically calculated as follows:
$W^{(\pm)}_{k_{0}\rightarrow k}$ $=$ $U_{\pm,k}^{(l){\dagger}}{U}_{\pm,k_{l-1}}^{(r)}\dots U_{\pm,k_{i}}^{(l){\dagger}}{U}_{\pm,{k}_{i-1}}^{(r)}\dots U_{\pm,{k}_{1}}^{(l){\dagger}}{U}_{\pm,{k}_{0}}^{(r)}$ (158)
$=$ $U_{\pm,k}^{(l){\dagger}}P_{\pm,k_{l-1}}\dots P_{\pm,{k}_{i}}\dots P_{\pm,{k}_{1}}U_{\pm,{k}_{0}}^{(r)},$
@@
Appendix E Non-Abelian dynamics in $k$ space for other initial states
Figure 4: (a) The identical to Fig. 1(a) of the main text, for convenience to analyze the subsequent figures. (b)–(i) The same settings as in Figs. 1(b)–1(i) of the main text, except for the initial state being instead located in the second-lowest band at $k=0$ in (a).
Figure 5: (a) The identical to Fig. 1(a) of the main text, for convenience to analyze the subsequent figures. (b)–(i) The same settings as in Figs. 1(b)–1(i) of the main text, except for the initial state being instead the equal superposition of Bloch states in the lowest two bands at $k=0$ in (a).
In this appendix, we demonstrate the non-Abelian dynamics $|\psi(t)\rangle=\hat{u}_{k(t)}^{(r){\dagger}}\alpha(t)|0\rangle$ in $k$ space in the same way as in Fig. 3 of the main text, but for two other initial states: (1) the Bloch state located in the second-lowest band at $k=0$, i.e., $\alpha(0)=(0,1,0,0)^{T}$, and (2) the equal superposition of Bloch states in the lowest two bands at $k=0$, i.e., $\alpha(0)=(1,1,0,0)^{T}/\sqrt{2}$. The corresponding figures are plotted in Figs. 4 and 5, respectively.